OpenStudy (anonymous):

Ln(x+2)-LN(4x+3)=LN(1/x)

OpenStudy (anonymous):

ln(u) - ln(v) = ln(u/v) so: ln(x+2) - ln(4x+3) = ln((x+2)/(4x+3)) = ln(1/x) => 1/x = (x+2)/(4x+3) now solve like a normal eq.

OpenStudy (anonymous):

Then times both sides by 4x+3?

OpenStudy (anonymous):

Or by x/1?

OpenStudy (anonymous):

both :D do it one at a time though, its easier

OpenStudy (anonymous):

you should get 4x+ 3 = x(x+2)